OpenStudy (wade123):

@phi last one(:

OpenStudy (wade123):

Write the equation of the function g(x)= f(x-3)+1 and f(x)=x^3+1

OpenStudy (wade123):

@phi

OpenStudy (phi):

it's easier if you rename the variables let f(a)= a^3 + 1 now what is f(x-3) ? which means replace "a" in the formula with (x-3) can you do that ?

OpenStudy (phi):

don't do anything crazy. just erase the "a" and put in (x-3) in f(a)= a^3 + 1